### v1

# Usman Ahmed,a Mikael P. Johansson,a,b Susi Lehtola,a and Dage Sundholm∗,a

Abstract
Hydrogen bonding is an important non-covalent interaction that plays a major role in molecular self-organization and supramolecular structures. It can be described accurately with ab initio quantum chemical wave function methods, which become computationally expensive for large molecular assemblies. Density functional theory (DFT) offers a better balance between accuracy and computational cost, and can be routinely applied to large systems. A large number of density functional approximations (DFAs) has been developed, but their accuracy depend on the application, necessitating benchmark studies to guide their selection for use in applications. Some of us have recently determined highly accurate hydrogen bonding energies of 14 quadruply hydrogen-bonded dimers by extrapolating coupled-cluster energies to the complete basis set limit as well as extrapolating electron correlation contributions with a continued-fraction approach [U. Ahmed et al., Phys. Chem. Chem. Phys., 2024, 26, 24470]. In this work, we study the reproduction of these bonding energies at the DFT level using 152 DFAs. The top ten functionals are composed of eight variants of the Berkeley functionals both with and without dispersion corrections, and two Minnesota 2011 functionals augmented with a further dispersion correction. We find the B97M-V functional with the non-local correlation functional replaced by an empirical D3BJ dispersion correction to be the best functional, while changes to the dispersion part in other Berkeley functionals lead to poorer performance in our study.
1 Introduction
Molecular self-assembly is a spontaneous process where disordered molecules organize into ordered structures without external influence.2, 3 The process leads to large supramolecular architectures, which are complexed through non-covalent interactions that guide the self-assembly. Understanding the complex self-assembly mechanism is necessary for the design of large supramolecular systems.
Hydrogen bonding is one of the most important non-covalent interactions governing the self-assembly due to its strength, selectivity, and directionality.4, 5 The significance of hydrogen bonding for stabilizing molecules and biomolecules is well known from, e. g., its role in the structure of deoxyribonucleic acid (DNA). The strength of hydrogen bonding depends strongly on the number of bonds, each hydrogen bond increasing the strength of the bonding. Molecules with four hydrogen bonds per molecule are major examples of multiple hydrogen bonding in self-organized systems. Self-complementarity enables identical molecules to interact and form larger and uniform supramolecular structures. Arrays of four hydrogen bonds can bind together in either a DDAA-AADD or a DADA-ADAD bonding motif, where D and A denote hydrogen bond donors and acceptors, respectively. These motifs are particularly consequential for designing stable assemblies with a given structure.
[…]
Accurate computational analysis of the strength and stability of supramolecular systems with hydrogen bonding is challenging, because the molecules are large. While ab initio correlated methods do offer an accurate description of hydrogen bonding, the rapidly increasing computational costs hamper computational studies of large molecules. Kohn–Sham density functional theory (DFT)21 has become the method of choice for studying such systems.22 DFT is widely used due to its cost effectiveness and availability in all modern quantum chemical software packages. Since the exact exchange-correlation functional is not known, more than 600 density functional approximations (DFAs) have so far been proposed and implemented.23 Since the accuracy of the DFAs varies depending on the application, it is not obvious which functional should be used for studying molecules with multiple hydrogen bonds.
The recent work by Montero de Hijes et al.24 on the density isobars of water and the melting temperature of ice based on machine learned density functional potential energy surfaces underlines the need for reinvestigating DFT models of hydrogen bonds. Their study found small changes in the semi-empirical dispersion correction in DFT calculations to result in large changes in the bulk properties of water and ice. As multiple hydrogen bonding carries a significant role for the structure of water, identifying functionals (which include dispersion corrections) that work especially well for modeling highly hydrogen bonded systems is useful to guide future studies of complex liquids and other systems where hydrogen-bonding interactions play a major role.
The accuracy of various DFAs for modeling hydrogen bonds has been evaluated in a number of works.25, 26, 27, 28, 29 More recent DFA benchmark studies of hydrogen-bonding strengths have employed the S22, S66 and S66x8 datasets.30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45 The pronounced relevance of hydrogen bonding in chemistry and their DFT modeling is underlined by the fact that hydrogen bond strengths have been included in the fitting of the Minnesota functionals since their first versions,29, 46, 47 for example.
In this work, we study how well various DFAs reproduce the hydrogen bonding in 14 quadruply hydrogen-bonded dimers. Half of the molecules exhibit the DDAA-AADD and the other half exhibits the DADA-ADAD hydrogen bonding motif. We consider 152 DFAs including the Hartree–Fock (HF) method. The employed dispersion correction schemes include Grimme’s empirical dispersion corrections (DFT-D348, DFT-D449) and a non-local (NL) correlation functional (VV10).50
Some of us have recently reported accurate reference values for the molecules studied in this work, which were calculated with domain-based local pair natural orbital coupled-cluster theory at the singles, doubles and perturbative triples DLPNO-CCSD(T) level extrapolated to the complete basis set limit.1 Ahmed et al.1 also extrapolated the reference energies towards the Schrödinger limit following Goodson51, producing the data that we will use as reference in this work.
This work is outlined as follows. The computational details are given in Section 2 along with the list of the studied DFAs. The obtained results are presented in Section 3: basis-set convergence is discussed in Section 3.1, and the results are analyzed in Section 3.2. The work is summarized and its main conclusions are drawn in Section 4.
2 Computational methods
The benchmark considers the 14 molecules of the study of Ahmed et al.1 that determined hydrogen-bonding energies for TPSSh-D3/def2-TZVPP optimized geometries at the DLPNO-CCSD(T) level of theory extrapolated to the complete basis set limit, as well as towards the Schrödinger limit using the continued-fraction (cf) scheme of Goodson51. We denote the reference level of theory as CCSD(T)-cf. The reference energies are taken from Ahmed et al.1 and are included in the ESI†along with the employed geometries and visualizations thereof.
[…]
The calculations were performed using the default (75,302) exchange-correlation grid in Psi4, which uses 75 radial grid points placed following Treutler and Ahlrichs55, and an angular Lebedev grid with 302 points.56 The VV10 non-local correlation functional50 was likewise evaluated on the corresponding (50,146) default grid. The resolution-of-the-identity approximation was used in all calculations with the universal Hartree–Fock auxiliary basis sets.57
The 75 base DFAs considered in this work, which have been listed in Table 1, comprise 14 semi-local generalized gradient approximation (GGA) functionals, 21 global hybrid (GH) GGA functionals, 13 range-separated (RS) GGA functionals, 13 semi-local meta-GGA (mGGA) functionals, 8 GH mGGA functionals, and 5 RS mGGA functionals, as well as HF (a GH functional without a semi-local exchange-correlation part). All of the density functionals were evaluated with Libxc,23 as usual in Psi4. Versions 6.0.0 and 6.2.2 of the library were employed. We note that some functionals such as B3LYP, HSE03, and HSE06 have various definitions in the literature,58 which is avoided through the use of the standard implementations in Libxc. Table 1 also shows the empirical (DFT-D348, DFT-D449) dispersion corrections or non-local correlation (VV1050) additionally considered with some of the functionals. Various combinations of these dispersion models (or the lack thereof) with the 75 base DFAs yield the 152 DFAs investigated in this study.
Table 1: The density functional approximations considered in this work. The three dispersion correction schemes are also indicated.
functional type D3 D4 NL ref. functional type D3 D4 NL ref.
B97-D3 GGA 59a 59 $\omega$B97 RS GGA 60
BLYP GGA 48a 49 61 62, 63, 64 $\omega$B97X RS GGA 49 60
BOP GGA 48a 65 $\omega$B97X-Db RS GGA 66
BP86 GGA 35a 67 62, 68 $\omega$B97X-V RS GGA 69c 70
CHACHIYO GGA 71, 72 $\omega$PBE RS GGA 48a 73, 74
KT1 GGA 75 CAM-B3LYP RS GGA 48a 76
KT2 GGA 75 CAMh-B3LYP RS GGA 77
KT3 GGA 78 CAM-QTP-00 RS GGA 79
N12 GGA 80a 81 CAM-QTP-01 RS GGA 82
PBE GGA 48a 49 67 83, 84 CAM-QTP-02 RS GGA 85
revPBE GGA 48a 49 61 86 HSE03 RS GGA 48a 49 87, 88
rPBE GGA 48a 49 89 HSE06 RS GGA 48a 49 87, 88, 90
PW91 GGA 91a 49 92, 93 N12-SX RS GGA 80a 94
XLYP GGA 48a 49 95 B97M-V mGGA 69c 96
HF GH 48a 49 M06-L mGGA 97
B1LYP GH GGA 48a 49 98 revM06-L mGGA 99
B3LYP GH GGA 48a 49 61 100 M11-L mGGA 80a 101
revB3LYP GH GGA 102 MN12-L mGGA 80a 103
B3PW91 GH GGA 48a 61 104 MN15-L mGGA 105
B3P86 GH GGA 35a 104 MVS mGGA 106, 107
B97 GH GGA 48c 108 SCAN mGGA 48a 49 109
B97-1 GH GGA 35a 110 rSCAN mGGA 111
B97-2 GH GGA 35a 112 r2SCAN mGGA 48a 113 114, 115
B97-3 GH GGA 116 TASK mGGA 117, 118
BHandH GH GGA 119 TPSS mGGA 48a 49 61 120, 121
BHandHLYP GH GGA 62, 119 revTPSS mGGA 48a 49 122 107, 123
BMK GH GGA 48a 124 M06 GH mGGA 31
HCTH120 GH GGA 48a 125 revM06 GH mGGA 126
HCTH407 GH GGA 48a 127 M06-2X GH mGGA 31
MPW1PW GH GGA 48a 49 128 M08-HX GH mGGA 129
O3LYP GH GGA 48a 49 130, 131 M08-SO GH mGGA 129
PBE0 GH GGA 48a 49 67 132, 133 MN15 GH mGGA 48a 134
revPBE0 GH GGA 48a 49 61 86, 83, 84 TPSSh GH mGGA 48a 49 27
QTP17 GH GGA 135 revTPSSh GH mGGA 48a 49 136
SOGGA11-X GH GGA 80a 137 $\omega$B97M-V RS mGGA 69c 138
X3LYP GH GGA 48a 49 95 M06-SX RS mGGA 139
M11 RS mGGA 48a 140
revM11 RS mGGA 141
MN12-SX RS mGGA 80a 94
  • a
    Empirical dispersion with Becke–Johnson142 3-body damping.
  • b
    The functional was parametrized with Grimme’s original dispersion correction, which is not available in Psi4, so the calculations with this functional do not include dispersion.
  • c
    Empirical dispersion with Becke–Johnson142 2-body damping.
3 Results
3.1 Basis-set convergence
We start the analysis by determining the basis-set convergence of the hydrogen bonding energies. We use the B97M-D3 functional and use the largest basis set considered (def2-QZVPPD) to determine reference values for this part of the study. The resulting basis-set truncation errors in the hydrogen-bonding energies are shown in Fig. 1 as a function of basis set.
Figure 1(a) shows the convergence for the basis sets without diffuse functions. The truncation errors decrease systematically, as the basis set is increased from split-valence with polarization to polarized valence triple- and quadruple-$\zeta$. The convergence is relatively slow with the cardinal number, and not all triple-$\zeta$ results are not yet converged to ±$\pm 0.5$ kJ/mol precision.
Figure 1(b) shows the analogous results including diffuse functions. The range of truncation errors for the polarized split-valence basis sets drops significantly from $\approx$ 0.8–7 kJ/mol with def2-SVP in Fig. 1(a) to $\approx$ $-1.5$–0.5 kJ/mol with def2-SVPD; this highlights the need to include diffuse functions for describing non-covalent bonding, as in the present hydrogen bonded systems. A convergence to within ±$\pm 0.5$ kJ/mol precision with respect to the basis set is already achieved with the def2-TZVPD basis set. It is thus clear from the data in Fig. 1 that the def2-QZVPPD basis set should afford benchmark accuracy hydrogen-bond energies suitable for the present study, and we will use the def2-QZVPPD basis set for the rest of the study.
(a) Convergence without diffuse functions
(b) Convergence with diffuse functions
Fig.  1: Basis-set convergence plot of the hydrogen-bonding energies, relative to B97M-D3/def2-QZVPPD values. The figures show an energy window of ±$\pm 0.5$ kJ/mol as horizontal dashed lines.
3.2 The best functionals
The functionals were ranked based on their mean absolute errors (MAE). The ranking is shown in Tables 2 and 3 together with the corresponding MAEs and maximum absolute errors (MAX). The lowest MAE is obtained with the B97M-D3 functional, which is the B97M-V functional where the VV10 nonlocal correlation functional (indicated by -V in the functional’s name) has been replaced with an empirical D3 dispersion correction; the original B97M-V functional is ranked 4th. The second-lowest MAE is achieved with the $\omega$B97M-V functional, which is a RS GGA functional again with VV10 dispersion. Replacing VV10 with the D3 dispersion correction drops the ranking to the 7th place.
Table 2: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the 76 highest ranked functionals. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The functionals are ordered in increasing MAE.
Rank Functional MAE MAX Rank Functional MAE MAX
1 B97M-D3 1.82 5.24 39 BLYP-D4 6.36 14.27
2 $\omega$B97M-V 2.06 5.45 40 TPSS-D3 6.37 12.98
3 M11-L-D3 2.13 5.22 41 M08-HX 6.43 10.81
4 B97M-V 2.17 6.71 42 $\omega$B97X-D4 6.67 11.55
5 $\omega$B97 2.21 5.86 43 revTPSS-NL 6.73 12.11
6 $\omega$B97X-V 2.43 6.09 44 r2SCAN-D4 7.07 13.55
7 $\omega$B97M-D3 2.48 5.24 45 TPSSh-D3 7.15 12.97
8 $\omega$B97X-D3 3.09 6.25 46 r2SCAN-D3 7.29 13.63
9 $\omega$B97X 3.11 7.98 47 HF-D4 7.31 14.14
10 M11-D3 3.17 7.27 48 PBE-D3 7.65 14.64
11 revTPSSh-D3 3.20 7.55 49 PW91-D3 7.70 14.23
12 SCAN 3.43 8.91 50 revM11 7.98 13.51
13 revTPSS-D3 3.48 9.21 51 TPSS-D4 7.99 15.50
14 revPBE-NL 3.64 8.94 52 B1LYP-D3 8.12 12.75
15 revTPSS-D4 3.65 9.64 53 M06-L 8.26 13.41
16 revTPSSh-D4 3.99 8.98 54 N12-SX 8.29 25.27
17 XLYP-D3 4.21 10.57 55 MN12-SX-D3 8.45 14.75
18 $\omega$PBE-D3 4.24 12.62 56 BMK-D3 8.49 14.50
19 rPBE-D4 4.30 10.82 57 CAM-QTP-01 8.54 20.80
20 B97-D3 4.45 12.54 58 BLYP-NL 8.66 13.94
21 BLYP-D3 4.58 11.96 59 TPSSh-D4 8.82 14.81
22 XLYP-D4 4.72 11.76 60 PBE-D4 9.26 16.42
23 BOP-D3 4.73 12.09 61 HSE03 9.35 29.92
24 HCTH407-D3 4.78 15.06 62 SCAN-D4 9.75 16.54
25 revPBE-D4 4.83 12.10 63 M06 9.85 16.87
26 HCTH120-D3 4.84 14.53 64 B3PW91-D3 9.92 16.30
27 rPBE-D3 5.03 16.16 65 M11 9.98 14.92
28 revPBE0-D3 5.05 8.89 66 B3LYP-D3 10.05 14.89
29 M06-2X 5.12 9.73 67 B1LYP-D4 10.09 14.61
30 revPBE-D3 5.37 15.66 68 MPW1PW-D3 10.14 16.95
31 revPBE0-NL 5.42 11.61 69 MN12-L-D3 10.25 17.57
32 CAM-QTP-00 5.53 15.23 70 PW91-D4 10.48 17.06
33 rSCAN 5.56 16.54 71 PBE-NL 10.65 16.64
34 B97-1-D3 5.59 10.42 72 MPW1PW-D4 10.74 17.73
35 MVS 5.66 13.53 73 HSE06 10.74 31.54
36 B97-2-D3 5.89 10.47 74 PBE0-D3 10.82 17.40
37 r2SCAN 5.93 15.69 75 SCAN-D3 10.88 17.64
38 revPBE0-D4 6.01 11.03 76 TPSS-NL 10.88 16.88
Table 3: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the functionals ranked from 77 to 152. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The functionals are ordered in increasing MAE.
Rank Functional MAE MAX Rank Functional MAE MAX
77 B3LYP-D4 11.05 15.90 115 BMK 19.84 31.71
78 BP86-D3 11.10 19.82 116 X3LYP 19.96 40.69
79 revM06 11.20 15.90 117 N12 20.22 48.75
80 X3LYP-D3 11.26 16.17 118 CAMh-B3LYP 20.44 41.43
81 HSE06-D3 11.35 17.78 119 $\omega$B97X-D 20.59 42.63
82 B3PW91-NL 11.60 18.99 120 MN12-L 20.79 27.11
83 X3LYP-D4 11.68 16.45 121 MPW1PW 21.09 47.74
84 CAM-B3LYP-D3 11.78 17.57 122 TPSSh 23.77 50.62
85 PBE0-D4 11.90 18.31 123 KT3 23.86 51.05
86 CAM-B3LYP 12.10 29.56 124 $\omega$PBE 23.91 48.60
87 M08-SO 12.13 19.50 125 revTPSSh 24.32 46.15
88 SOGGA11-X-D3 12.13 17.82 126 TPSS 25.91 53.91
89 B3LYP-NL 12.18 17.31 127 B3LYP 26.54 50.90
90 PBE0-NL 12.22 19.09 128 revTPSS 26.68 49.15
91 HSE03-D3 12.53 19.42 129 BP86 27.33 58.38
92 MN15-D3 12.64 18.31 130 B1LYP 28.11 51.93
93 PW91 12.80 35.00 131 B97-0 28.81 53.34
94 PBE0 13.04 34.52 132 O3LYP-D4 29.21 51.57
95 HSE06-D4 13.12 19.58 133 O3LYP-D3 29.77 52.14
96 MN15 13.33 19.31 134 B3PW91 29.82 61.28
97 MN15-L 13.33 19.31 135 QTP-17 30.03 47.11
98 CAM-QTP-02 13.66 26.81 136 B97-3 30.95 53.53
99 M06-SX 13.71 19.06 137 KT1 31.33 39.56
100 B3P86-D3 14.21 20.56 138 B97-2 33.77 64.10
101 HSE03-D4 14.26 20.90 139 HCTH120 36.29 65.40
102 BHANDHLYP 15.21 33.09 140 revPBE0 37.87 69.50
103 SOGGA11-X 15.59 31.93 141 XLYP 38.30 65.73
104 M11-L 15.69 26.51 142 BLYP 39.54 69.55
105 BP86-NL 16.09 23.11 143 KT2 39.62 49.78
106 N12-SX-D3 16.34 21.96 144 BHANDH 44.53 61.65
107 HF-D3 16.63 25.87 145 HF 46.07 71.35
108 PBE 17.27 40.79 146 HCTH407 47.72 79.14
109 revM06-L 17.67 26.51 147 revPBE 50.92 88.16
110 B3P86 17.84 43.49 148 rPBE 51.03 85.81
111 B97-1 17.96 36.89 149 O3LYP 61.36 102.66
112 N12-D3 18.19 25.07 150 BOP 63.69 105.40
113 revB3LYP 19.61 41.55 151 CHACHIYO 63.76 111.80
114 MN12-SX 19.79 25.07 152 TASK 148.78 216.80
The Minnesota M11-L-D3 holds 3rd place on the list with the added D3 dispersion corrections, while the original functional parametrized without the empirical dispersion term (M11-L) holds the 104th place.
The functionals ranked $5^{\mathrm{th}}$ to $9^{\mathrm{th}}$ are all Berkeley $\omega$B97 (RS GGA) functionals with different dispersion correction terms. $\omega$B97X-D3 is the $\omega$B97X-V70 functional with a D3 Becke–Johnson142, 48 dispersion correction instead of VV1069, and it should not to be confused with the $\omega$B97X-D3 functional of Lin et al.143 which was not studied in this work.
The Minnesota M11 functional with an additional D3 correction is on the 10th place, while the original functional without the dispersion term ranks 65th.
The $\omega$B97X-D4 functional is ranked 42nd, which is 34 places below the rank of $\omega$B97X-D3. This is likely because $\omega$B97X-D4 is based on a different base functional than $\omega$B97X-D3: it is built on top of an older functional, $\omega$B97X instead of $\omega$B97X-V.49 While a dispersion correction was included in the fitting of the $\omega$B97X-V functional through the VV10 functional, $\omega$B97X was fit without dispersion terms.
Various mGGA functionals and one revised GGA functional are ranked 11th to 16th.
Five of the six studied Berkeley functionals are ranked among the 10 best performing DFAs. mGGA and global hybrid (GH) GGA functionals also appear among the 20 most accurate functionals for hydrogen-bonding energies. The best GGA functionals are ranked 17th (XLYP-D3), 19th (rPBE-D4), 21st (BLYP-D3) and 22nd (XLYP-D4).
Fig.  2: Violin plots showing the error distributions in the hydrogen-bonding energies (in kJ/mol) for the investigated DFAs. The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1
The error distributions for the hydrogen-bonding energies of the 36 top functionals are shown as violin plots in Fig. 2, while violin plots for the remaining functionals are included in the supporting information (SI). The average errors in the hydrogen-bonding energies for the top functionals are close to zero, except for M11-D3, whose distribution is almost completely on the negative side, implying that it systematically underestimates the hydrogen bonding energies. Similar plots are shown for the functionals ranked 13th–24th in Fig. 2, where the distributions are broader, and the average is larger than zero for many functionals, which implies that the hydrogen-bonding energies are overestimated. The error distributions for the functionals ranked between 25th and 36th shown in Fig. 2 are broader and more scattered than for the higher-ranked functionals. The violin plots for the rest of the functionals are given in the ESI†.
3.3 Meta-GGA functionals
The best performing mGGA functionals are B97M-D3 (1st), M11-L-D3 (3rd) and B97M-V (4th). SCAN and revTPSS-D3 are mGGA functionals that also perform well, holding rank 12 and 13, respectively. SCAN exhibits worse accuracy when paired with the D3 (ranked 75th) or D4 (ranked 62th) dispersion, suggesting overbinding. In contrast, the revTPSS functional improves its performance significantly by adding the empirical dispersion correction term to the functional, the rank increasing from 125th to 13th (D3) or 15th (D4), which suggests that the hydrogen bonds are too weak at the revTPSS level, which is rectified by the dispersion correction. The rest of the mGGA functionals among the first third part of the ranking list includes rSCAN (33rd), MVS (35th) and r2SCAN (37th). However, by considering the dispersion interaction using the D3 correction term, the performance of r2SCAN drops from 37th to 46th and to 44th with D4. Common mGGA functionals such as M06-L (53rd) and MN15-L (97th) are in the second third part of the ranking list, whereas M11-L without dispersion (104th), revM06-L (109th), MN12-L (120th), TPSS (126th), revTPSS (128th) and TASK (152nd) are in last third of the ranking list. The performance of the TPSS, revTPSS, MN12-L, and MN12-L functionals improves significantly when adding the empirical dispersion term. The accuracy is in a few cases improved by using D4 instead of D3, whereas using the VV10 dispersion correction leads to a lower accuracy than obtained with the D3 correction.
3.4 Range-separated hybrid functionals
The Berkeley range-separated functionals perform well. The $\omega$B97M-V is the best RS mGGA functional in the 2nd place and the Minnesota M11 functional combined with the D3 correlation term is ranked 10th. The best of the rest of the RS mGGA functionals is revM11, which is ranked 50th. Three of the five RS GGA functionals from Berkeley are ranked among the 10 best functionals, while adding a dispersion correction to $\omega$B97X leads to worse performace and 42nd rank. The $\omega$B97X-D functional performs poorly in our study (119th), but this is likely caused by lack of dispersion in our calculations with this functional; see Table 1.
The best RS GGA functional from Gainesville (QTP) is 32nd, whereas the rest of the RS GGA functionals are ranked between 54th and 135th. The short-range-only HSE03 and HSE06 functionals rank 61st and 73rd, respectively. Adding the dispersion D3 term to CAM-B3LYP has a very small effect on its MAE, whereas D3 improves its performance by decreasing the MAX value by 40%.
3.5 Global hybrid functionals
We have investigated the performance of 21 global hybrid GGA functionals and added the D3 dispersion terms to 16 of them. Their rank without the D3 term is in general 100th and larger. The only exceptions is PBE0 (94th). GH GGA functionals perform better when adding the D3 dispersion term. The best GH GGA functionals with the D3 correction are B97 (20th), HCTH407 (24th), HCTH120 (26th), revPBE0 (28th), B97-1 (34th) and B97-2 (36th). The ranking of the other GH GGA-D3 functionals is 52nd and larger. The use of D4 or NL dispersion corrections leads in general to a slightly less accurate functional than with D3.
The best global hybrid mGGA functional without additional dispersion term is M06-2X in the 29th place. The M08-HX functional is ranked 41st, whereas the rest of the GH mGGA functionals is ranked between 63rd and 125th. The performance of the TPSSh functional improves from 122nd to 45th (D3) and 59th (D4) when adding the empirical dispersion term to the functional. Adding an empirical dispersion correction to the revTPSSh functional improves its performance significantly: the ranking of revTPSSh increases from 125th to 11th (revTPSSh-D3) or 16th (revTPSSh-D4).
3.6 GGA functionals
The revPBE-NL and XLYP-D3 functionals are the best performing functionals among the GGAs. Their performance drops from the 14th place to 124th, and 17th rank to 141th without dispersion corrections, respectively. The rPBE-D4 and BLYP-D3 functionals are the next ones on the GGA ranking list ranked 19th and 21st, respectively. The BOP functional places 23rd. Other GGA functionals among the first 50 on the ranking list are PBE-D3 and PW91-D3, which are ranked 48th and 49th, respectively. The rest of the GGA functionals including dispersion effects appear among the second third on the ranking lists. The ranking of HF is 145th, which improves to 46th and 107th with D3 and D4 dispersion corrections, respectively.
4 Summary and conclusions
We calculated the hydrogen-bonding strength for 14 molecules using 152 density functional approximations (DFAs). The accuracy of the calculated bond strengths with various DFAs was assessed by comparison to the literature CCSD(T)-cf energies of Ahmed et al.1, which they extrapolated to the complete basis set limit, as well as to the Schrödinger limit using the continued-fraction methodology of Goodson51.
We began the work by validating the numerical reliability of our computational method by assessing the convergence of the hydrogen-bonding strengths across the Karlsruhe basis set family for the B97M-D3 functional. Taking the def2-QZVPPD values as reference, we observed that diffuse functions are important to obtain converged interaction energies, and we therefore recommend their use in modeling hydrogen bonds. As the def2-TZVPD and def2-TZVPPD values are within 0.5 kJ/mol of the def2-QZVPPD data, we can safely assume that the quadruple-$\zeta$ reference calculations are indeed of benchmark quality. The def2-QZVPPD basis set is often used for benchmarking in the literature, and has been employed in the fitting of the $\omega$B97M-V functional,138 for example.
Having validated the numerical methodology, we continued with the benchmark. We found the top of the list to be dominated by the Berkeley functionals, either with or without dispersion. The most accurate functional for hydrogen-bonding strengths was found to be B97M-D3, which is the B97M-V functional whose VV10 dispersion correction has been replaced by the empirical D3 dispersion correction. We found the accuracy of mGGA functionals to be surprisingly sensitive to the description of the dispersion energy, tentatively supporting the recent findings of Montero de Hijes et al.24. Other functionals that accurately predicted hydrogen-bonding strengths included the range-separated $\omega$B97M-V mGGA functional, the $\omega$B97, $\omega$B97X, $\omega$B97X-V range-separated GGA functionals, as well as the semi-local M11-L-D3 mGGA and the range-separated M11-D3 mGGA functional.
Acknowledgement
This work has been supported by the Academy of Finland through project numbers 314821, 340583, 350282, and 353749. It has also been financially supported by the Magnus Ehrnrooth Foundation, the Oskar Öflund Foundation, and by the Swedish Cultural Foundation in Finland. We acknowledge computational resources from CSC - IT Center for Science, Finland.
Notes and references
  • Ahmed et al. 2024 U. Ahmed, C. D. Daub, D. Sundholm and M. P. Johansson, Attractive acceptor–acceptor interactions in self-complementary quadruple hydrogen bonds for molecular self-assembly, Phys. Chem. Chem. Phys., 2024, 26, 24470–24476, DOI: 10.1039/D4CP02361G.
  • Whitesides et al. 1991 G. M. Whitesides, J. P. Mathias and C. T. Seto, Molecular Self-Assembly and Nanochemistry: a Chemical Strategy for the Synthesis of Nanostructures, Science, 1991, 254, 1312–1319, DOI: 10.1126/science.1962191.
  • Pochan and Scherman 2021 D. Pochan and O. Scherman, Introduction: Molecular Self-Assembly, Chem. Rev., 2021, 121, 13699–13700, DOI: 10.1021/acs.chemrev.1c00884.
  • Cooke and Rotello 2002 G. Cooke and V. M. Rotello, Methods of modulating hydrogen bonded interactions in synthetic host–guest systems, Chem. Soc. Rev., 2002, 31, 275–286, DOI: 10.1039/B103906G.
  • Buckingham et al. 2008 A. Buckingham, J. Del Bene and S. McDowell, The hydrogen bond, Chem. Phys. Lett., 2008, 463, 1–10, DOI: 10.1016/j.cplett.2008.06.060.
  • Beijer et al. 1998 F. H. Beijer, R. P. Sijbesma, H. Kooijman, A. L. Spek and E. W. Meijer, Strong Dimerization of Ureidopyrimidones via Quadruple Hydrogen Bonding, J. Am. Chem. Soc., 1998, 120, 6761–6769, DOI: 10.1021/ja974112a.
  • Lafitte et al. 2005 V. G. H. Lafitte, A. E. Aliev, H. C. Hailes, K. Bala and P. Golding, Ureidopyrimidinones Incorporating a Functionalizable p-Aminophenyl Electron-Donating Group at C-6, J. Org. Chem., 2005, 70, 2701–2707, DOI: 10.1021/jo048223l.
  • Corbin and Zimmerman 2000 P. S. Corbin and S. C. Zimmerman, Complexation-Induced Unfolding of Heterocyclic Ureas: A Hydrogen-Bonded Sheetlike Heterodimer, J. Am. Chem. Soc., 2000, 122, 3779–3780, DOI: 10.1021/ja992830m.
  • Moriuchi et al. 2002 T. Moriuchi, T. Tamura and T. Hirao, Self-Assembly of Dipeptidyl Ureas: A New Class of Hydrogen-Bonded Molecular Duplexes, J. Am. Chem. Soc., 2002, 124, 9356–9357, DOI: 10.1021/ja020098c.
  • Sanford et al. 2004 A. R. Sanford, K. Yamato, X. Yang, L. Yuan, Y. Han and B. Gong, Well-defined secondary structures, Eur. J. Biochem., 2004, 271, 1416–1425, DOI: 10.1111/j.1432-1033.2004.04062.x.
  • Zhao et al. 2003 X. Zhao, X.-Z. Wang, X.-K. Jiang, Y.-Q. Chen, Z.-T. Li and G.-J. Chen, Hydrazide-Based Quadruply Hydrogen-Bonded Heterodimers. Structure, Assembling Selectivity, and Supramolecular Substitution, J. Am. Chem. Soc., 2003, 125, 15128–15139, DOI: 10.1021/ja037312x.
  • Baruah et al. 2005 P. K. Baruah, R. Gonnade, U. D. Phalgune and G. J. Sanjayan, Self-Assembly with Degenerate Prototropy, J. Org. Chem., 2005, 70, 6461–6467, DOI: 10.1021/jo0508705.
  • Otte et al. 2013 P. Otte, J. Taubitz and U. Lüning, Lipophilicity Enhancing Substituents for ADDA Recognition Domains of DAAD·ADDA Heterodimers with Quadruple Hydrogen Bonds, Eur. J. Org. Chem., 2013, 2013, 2130–2139, DOI: /10.1002/ejoc.201201450.
  • Ośmiałowski et al. 2012 B. Ośmiałowski, E. Kolehmainen, S. Ikonen, A. Valkonen, A. Kwiatkowski, I. Grela and E. Haapaniemi, 2-Acylamino- and 2,4-Bis(acylamino)pyrimidines as Supramolecular Synthons Analyzed by Multiple Noncovalent Interactions. DFT, X-ray Diffraction, and NMR Spectral Studies, J. Org. Chem., 2012, 77, 9609–9619, DOI: 10.1021/jo301643z.
  • Lafitte et al. 2011 V. G. H. Lafitte, A. E. Aliev, E. Greco, K. Bala, P. Golding and H. C. Hailes, Quadruple hydrogen bonded cytosine modules: N-1 functionalised arrays, New J. Chem., 2011, 35, 1522–1527, DOI: 10.1039/C1NJ20162J.
  • Greco et al. 2010 E. Greco, A. E. Aliev, V. G. H. Lafitte, K. Bala, D. Duncan, L. Pilon, P. Golding and H. C. Hailes, Cytosine modules in quadruple hydrogen bonded arrays, New J. Chem., 2010, 34, 2634–2642, DOI: 10.1039/C0NJ00197J.
  • Martin et al. 2006 A. M. Martin, R. S. Butler, I. Ghiviriga, R. E. Giessert, K. A. Abboud and R. K. Castellano, Self-complementary purines by quadruple hydrogen bonding, Chem. Comm., 2006, 4413–4415, DOI: 10.1039/B610239E.
  • Kheria et al. 2017 S. Kheria, S. Rayavarapu, A. S. Kotmale, R. G. Gonnade and G. J. Sanjayan, Triazine-Based Highly Stable AADD-Type Self-Complementary Quadruple Hydrogen-Bonded Systems Devoid of Prototropy, Chem. Eur. J., 2017, 23, 783–787, DOI: 10.1002/chem.201605208.
  • Baruah and Khan 2013 P. K. Baruah and S. Khan, Self-complementary quadruple hydrogen bonding motifs: from design to function, RSC Adv., 2013, 3, 21202–21217, DOI: 10.1039/C3RA43814G.
  • Ahmed et al. 2024 U. Ahmed, D. Sundholm and M. P. Johansson, The effect of hydrogen bonding on the $\pi$ depletion and the $\pi$$\pi$ stacking interaction, Phys. Chem. Chem. Phys., 2024, 26, 27431–27438, DOI: 10.1039/D4CP02889A.
  • Kohn and Sham 1965 W. Kohn and L. J. Sham, Self-Consistent Equations Including Exchange and Correlation Effects, Phys. Rev., 1965, 140, A1133–A1138, DOI: 10.1103/PhysRev.140.A1133.
  • Becke 2014 A. D. Becke, Perspective: Fifty years of density-functional theory in chemical physics, J. Chem. Phys., 2014, 140, 18A301, DOI: 10.1063/1.4869598.
  • Lehtola et al. 2018 S. Lehtola, C. Steigemann, M. J. T. Oliveira and M. A. L. Marques, Recent developments in LIBXC—a comprehensive library of functionals for density functional theory, SoftwareX, 2018, 7, 1–5, DOI: 10.1016/j.softx.2017.11.002.
  • Montero de Hijes et al. 2024 P. Montero de Hijes, C. Dellago, R. Jinnouchi and G. Kresse, Density isobar of water and melting temperature of ice: Assessing common density functionals, J. Chem. Phys., 2024, 161, 131102, DOI: 10.1063/5.0227514.
  • Tuma et al. 1999 C. Tuma, A. D. Boese and N. C. Handy, Predicting the binding energies of H-bonded complexes: A comparative DFT study, Phys. Chem. Chem. Phys., 1999, 1, 3939–3947, DOI: 10.1039/A904357H.
  • Rabuck and Scuseria 2000 A. Rabuck and G. Scuseria, Performance of recently developed kinetic energy density functionals for the calculation of hydrogen binding strengths and hydrogen-bonded structures, Theo. Chem. Acc., 2000, 104, 439–444, DOI: 10.1007/s002140000163.
  • Staroverov et al. 2003 V. N. Staroverov, G. E. Scuseria, J. Tao and J. P. Perdew, Comparative assessment of a new nonempirical density functional: Molecules and hydrogen-bonded complexes, J. Chem. Phys., 2003, 119, 12129–12137, DOI: 10.1063/1.1626543.
  • Xu and Goddard 2004 X. Xu and W. A. Goddard, Bonding Properties of the Water Dimer: A Comparative Study of Density Functional Theories, J. Phys. Chem. A, 2004, 108, 2305–2313, DOI: 10.1021/jp035869t.
  • Zhao and Truhlar 2005 Y. Zhao and D. G. Truhlar, Benchmark Databases for Nonbonded Interactions and Their Use To Test Density Functional Theory, J. Chem. Theory Comput., 2005, 1, 415–432, DOI: 10.1021/ct049851d.
  • Zhao and Truhlar 2007 Y. Zhao and D. G. Truhlar, Density Functionals for Noncovalent Interaction Energies of Biological Importance, J. Chem. Theory Comput., 2007, 3, 289–300, DOI: 10.1021/ct6002719.
  • Zhao and Truhlar 2008 Y. Zhao and D. G. Truhlar, The M06 suite of density functionals for main group thermochemistry, thermochemical kinetics, noncovalent interactions, excited states, and transition elements: two new functionals and systematic testing of four M06-class functionals and 12 other function, Theor. Chem. Acc., 2008, 120, 215–241, DOI: 10.1007/s00214-007-0310-x.
  • Jurečka et al. 2006 P. Jurečka, J. Šponer, J. Černý and P. Hobza, Benchmark database of accurate (MP2 and CCSD(T) complete basis set limit) interaction energies of small model complexes, DNA base pairs, and amino acid pairs, Phys. Chem. Chem. Phys., 2006, 8, 1985–1993, DOI: 10.1039/B600027D.
  • Zhao and Truhlar 2008 Y. Zhao and D. G. Truhlar, Density Functionals with Broad Applicability in Chemistry, Acc. Chem. Res., 2008, 41, 157–167, DOI: 10.1021/ar700111a.
  • Goerigk et al. 2011 L. Goerigk, H. Kruse and S. Grimme, Benchmarking Density Functional Methods against the S66 and S66x8 Datasets for Non-Covalent Interactions, ChemPhysChem, 2011, 12, 3421–3433, DOI: 10.1002/cphc.201100826.
  • Goerigk et al. 2017 L. Goerigk, A. Hansen, C. Bauer, S. Ehrlich, A. Najibi and S. Grimme, A look at the density functional theory zoo with the advanced GMTKN55 database for general main group thermochemistry, kinetics and noncovalent interactions, Phys. Chem. Chem. Phys., 2017, 19, 32184–32215, DOI: 10.1039/C7CP04913G.
  • Burns et al. 2011 L. A. Burns, Á. Mayagoitia-Vázquez, B. G. Sumpter and C. D. Sherrill, Density-functional approaches to noncovalent interactions: A comparison of dispersion corrections (DFT-D), exchange-hole dipole moment (XDM) theory, and specialized functionals, J. Chem. Phys., 2011, 134, 084107, DOI: 10.1063/1.3545971.
  • Aragó et al. 2013 J. Aragó, E. Ortí and J. C. Sancho-García, Nonlocal van der Waals Approach Merged with Double-Hybrid Density Functionals: Toward the Accurate Treatment of Noncovalent Interactions, J. Chem. Theory Comput., 2013, 9, 3437–3443, DOI: 10.1021/ct4003527.
  • DiLabio et al. 2013 G. A. DiLabio, E. R. Johnson and A. Otero-de-la Roza, Performance of conventional and dispersion-corrected density-functional theory methods for hydrogen bonding interaction energies, Phys. Chem. Chem. Phys., 2013, 15, 12821–12828, DOI: 10.1039/C3CP51559A.
  • Roch and Baldridge 2017 L. M. Roch and K. K. Baldridge, Dispersion-Corrected Spin-Component-Scaled Double-Hybrid Density Functional Theory: Implementation and Performance for Non-covalent Interactions, J. Chem. Theory Comput., 2017, 13, 2650–2666, DOI: 10.1021/acs.jctc.7b00220.
  • Remya and Suresh 2013 K. Remya and C. H. Suresh, Which density functional is close to CCSD accuracy to describe geometry and interaction energy of small noncovalent dimers? A benchmark study using Gaussian09, J. Comp. Chem., 2013, 34, 1341–1353, DOI: 10.1002/jcc.23263.
  • Jurečka et al. 2007 P. Jurečka, J. Černý, P. Hobza and D. R. Salahub, Density functional theory augmented with an empirical dispersion term. Interaction energies and geometries of 80 noncovalent complexes compared with ab initio quantum mechanics calculations, J. Comp. Chem., 2007, 28, 555–569, DOI: 10.1002/jcc.20570.
  • Brauer et al. 2016 B. Brauer, M. K. Kesharwani, S. Kozuch and J. M. L. Martin, The S66x8 benchmark for noncovalent interactions revisited: explicitly correlated ab initio methods and density functional theory, Phys. Chem. Chem. Phys., 2016, 18, 20905–20925, DOI: 10.1039/C6CP00688D.
  • Santra et al. 2022 G. Santra, E. Semidalas, N. Mehta, A. Karton and J. M. L. Martin, S66x8 noncovalent interactions revisited: new benchmark and performance of composite localized coupled-cluster methods, Phys. Chem. Chem. Phys., 2022, 24, 25555–25570, DOI: 10.1039/D2CP03938A.
  • Řezáč 2020 J. Řezáč, Non-Covalent Interactions Atlas Benchmark Data Sets: Hydrogen Bonding, J. Chem. Theory Comput., 2020, 16, 2355–2368, DOI: 10.1021/acs.jctc.9b01265.
  • Řezáč 2020 J. Řezáč, Non-Covalent Interactions Atlas Benchmark Data Sets 2: Hydrogen Bonding in an Extended Chemical Space, J. Chem. Theory Comput., 2020, 16, 6305–6316, DOI: 10.1021/acs.jctc.0c00715.
  • Zhao et al. 2005 Y. Zhao, N. E. Schultz and D. G. Truhlar, Exchange-correlation functional with broad accuracy for metallic and nonmetallic compounds, kinetics, and noncovalent interactions, J. Chem. Phys., 2005, 123, 161103, DOI: 10.1063/1.2126975.
  • Zhao et al. 2006 Y. Zhao, N. E. Schultz and D. G. Truhlar, Design of Density Functionals by Combining the Method of Constraint Satisfaction with Parametrization for Thermochemistry, Thermochemical Kinetics, and Noncovalent Interactions, J. Chem. Theory Comput., 2006, 2, 364–382, DOI: 10.1021/ct0502763.
  • Grimme et al. 2011 S. Grimme, S. Ehrlich and L. Goerigk, Effect of the damping function in dispersion corrected density functional theory, J. Comput. Chem., 2011, 32, 1456–1465, DOI: 10.1002/jcc.21759.
  • Caldeweyher et al. 2019 E. Caldeweyher, S. Ehlert, A. Hansen, H. Neugebauer, S. Spicher, C. Bannwarth and S. Grimme, A generally applicable atomic-charge dependent London dispersion correction, J. Chem. Phys., 2019, 150, 154122, DOI: 10.1063/1.5090222.
  • Vydrov and Van Voorhis 2010 O. A. Vydrov and T. Van Voorhis, Nonlocal van der Waals density functional: the simpler the better., J. Chem. Phys., 2010, 133, 244103, DOI: 10.1063/1.3521275.
  • Goodson 2002 D. Z. Goodson, Extrapolating the coupled-cluster sequence toward the full configuration-interaction limit, J. Chem. Phys., 2002, 116, 6948–6956, DOI: 10.1063/1.1462620.
  • Smith et al. 2020 D. G. A. Smith, L. A. Burns, A. C. Simmonett, R. M. Parrish, M. C. Schieber, R. Galvelis, P. Kraus, H. Kruse, R. Di Remigio, A. Alenaizan, A. M. James, S. Lehtola, J. P. Misiewicz, M. Scheurer, R. A. Shaw, J. B. Schriber, Y. Xie, Z. L. Glick, D. A. Sirianni, J. S. O’Brien, J. M. Waldrop, A. Kumar, E. G. Hohenstein, B. P. Pritchard, B. R. Brooks, H. F. Schaefer, A. Y. Sokolov, K. Patkowski, A. E. DePrince, U. Bozkaya, R. A. King, F. A. Evangelista, J. M. Turney, T. D. Crawford and C. D. Sherrill, Psi4 1.4: Open-source software for high-throughput quantum chemistry, J. Chem. Phys., 2020, 152, 184108, DOI: 10.1063/5.0006002.
  • Weigend and Ahlrichs 2005 F. Weigend and R. Ahlrichs, Balanced basis sets of split valence, triple zeta valence and quadruple zeta valence quality for H to Rn: Design and assessment of accuracy, Phys. Chem. Chem. Phys., 2005, 7, 3297–3305.
  • Rappoport and Furche 2010 D. Rappoport and F. Furche, Property-optimized Gaussian basis sets for molecular response calculations, J. Chem. Phys., 2010, 133, 134105, DOI: 10.1063/1.3484283.
  • Treutler and Ahlrichs 1995 O. Treutler and R. Ahlrichs, Efficient molecular numerical integration schemes, J. Chem. Phys., 1995, 102, 346–354, DOI: 10.1063/1.469408.
  • Lebedev 1995 V. I. Lebedev, A quadrature formula for the sphere of 59th algebraic order of accuracy, Russ. Acad. Sci. Dokl. Math., 1995, 50, 283–286.
  • Weigend 2008 F. Weigend, Hartree–Fock exchange fitting basis sets for H to Rn, J. Comput. Chem., 2008, 29, 167–175, DOI: 10.1002/jcc.20702.
  • Lehtola and Marques 2023 S. Lehtola and M. A. L. Marques, Reproducibility of density functional approximations: How new functionals should be reported, J. Chem. Phys., 2023, 159, 114116, DOI: 10.1063/5.0167763.
  • Grimme 2006 S. Grimme, Semiempirical GGA-type density functional constructed with a long-range dispersion correction, J. Comput. Chem., 2006, 27, 1787–1799, DOI: 10.1002/jcc.20495.
  • Chai and Head-Gordon 2008 J.-D. Chai and M. Head-Gordon, Systematic optimization of long-range corrected hybrid density functionals, J. Chem. Phys., 2008, 128, 084106, DOI: 10.1063/1.2834918.
  • Hujo and Grimme 2011 W. Hujo and S. Grimme, Performance of the van der Waals Density Functional VV10 and (hybrid)GGA Variants for Thermochemistry and Noncovalent Interactions, J. Chem. Theory Comput., 2011, 7, 3866–3871, DOI: 10.1021/ct200644w.
  • Becke 1988 A. D. Becke, Density-functional exchange-energy approximation with correct asymptotic behavior, Phys. Rev. A, 1988, 38, 3098–3100, DOI: 10.1103/PhysRevA.38.3098.
  • Lee et al. 1988 C. Lee, W. Yang and R. G. Parr, Development of the Colle–Salvetti correlation-energy formula into a functional of the electron density, Phys. Rev. B, 1988, 37, 785–789, DOI: 10.1103/PhysRevB.37.785.
  • Miehlich et al. 1989 B. Miehlich, A. Savin, H. Stoll and H. Preuss, Results obtained with the correlation energy density functionals of Becke and Lee, Yang and Parr, Chem. Phys. Lett., 1989, 157, 200–206, DOI: 10.1016/0009-2614(89)87234-3.
  • Tsuneda et al. 1999 T. Tsuneda, T. Suzumura and K. Hirao, A new one-parameter progressive Colle–Salvetti-type correlation functional, J. Chem. Phys., 1999, 110, 10664–10678, DOI: 10.1063/1.479012.
  • Chai and Head-Gordon 2008 J.-D. Chai and M. Head-Gordon, Long-range corrected hybrid density functionals with damped atom-atom dispersion corrections, Phys. Chem. Chem. Phys., 2008, 10, 6615–6620, DOI: 10.1039/b810189b.
  • Kesharwani et al. 2016 M. K. Kesharwani, A. Karton and J. M. L. Martin, Benchmark ab Initio Conformational Energies for the Proteinogenic Amino Acids through Explicitly Correlated Methods. Assessment of Density Functional Methods, J. Chem. Theory Comput., 2016, 12, 444–454, DOI: 10.1021/acs.jctc.5b01066.
  • Perdew 1986 J. P. Perdew, Density-functional approximation for the correlation energy of the inhomogeneous electron gas, Phys. Rev. B, 1986, 33, 8822–8824, DOI: 10.1103/PhysRevB.33.8822.
  • Najibi and Goerigk 2018 A. Najibi and L. Goerigk, The Nonlocal Kernel in van der Waals Density Functionals as an Additive Correction: An Extensive Analysis with Special Emphasis on the B97M-V and $\omega$B97M-V Approaches, J. Chem. Theory Comput., 2018, 14, 5725–5738, DOI: 10.1021/acs.jctc.8b00842.
  • Mardirossian and Head-Gordon 2014 N. Mardirossian and M. Head-Gordon, $\omega$B97X-V: A 10-parameter, range-separated hybrid, generalized gradient approximation density functional with nonlocal correlation, designed by a survival-of-the-fittest strategy, Phys. Chem. Chem. Phys., 2014, 16, 9904–9924, DOI: 10.1039/c3cp54374a.
  • Chachiyo and Chachiyo 2020 T. Chachiyo and H. Chachiyo, Simple and Accurate Exchange Energy for Density Functional Theory, Molecules, 2020, 25, 3485, DOI: 10.3390/molecules25153485.
  • Chachiyo and Chachiyo 2020 T. Chachiyo and H. Chachiyo, Understanding electron correlation energy through density functional theory, Comput. Theor. Chem., 2020, 1172, 112669, DOI: 10.1016/j.comptc.2019.112669.
  • Henderson et al. 2008 T. M. Henderson, B. G. Janesko and G. E. Scuseria, Generalized gradient approximation model exchange holes for range-separated hybrids., J. Chem. Phys., 2008, 128, 194105, DOI: 10.1063/1.2921797.
  • Weintraub et al. 2009 E. Weintraub, T. M. Henderson and G. E. Scuseria, Long-Range-Corrected Hybrids Based on a New Model Exchange Hole, J. Chem. Theory Comput., 2009, 5, 754–762, DOI: 10.1021/ct800530u.
  • Keal and Tozer 2003 T. W. Keal and D. J. Tozer, The exchange-correlation potential in Kohn–Sham nuclear magnetic resonance shielding calculations, J. Chem. Phys., 2003, 119, 3015–3024, DOI: 10.1063/1.1590634.
  • Yanai et al. 2004 T. Yanai, D. P. Tew and N. C. Handy, A new hybrid exchange–correlation functional using the Coulomb-attenuating method (CAM-B3LYP), Chem. Phys. Lett., 2004, 393, 51–57, DOI: 10.1016/j.cplett.2004.06.011.
  • Shao et al. 2020 Y. Shao, Y. Mei, D. Sundholm and V. R. I. Kaila, Benchmarking the Performance of Time-Dependent Density Functional Theory Methods on Biochromophores, J. Chem. Theory Comput., 2020, 16, 587–600, DOI: 10.1021/acs.jctc.9b00823.
  • Keal and Tozer 2004 T. W. Keal and D. J. Tozer, A semiempirical generalized gradient approximation exchange-correlation functional, J. Chem. Phys., 2004, 121, 5654–5660, DOI: 10.1063/1.1784777.
  • Verma and Bartlett 2014 P. Verma and R. J. Bartlett, Increasing the applicability of density functional theory. IV. Consequences of ionization-potential improved exchange-correlation potentials, J. Chem. Phys., 2014, 140, 18A534, DOI: 10.1063/1.4871409.
  • Goerigk 2015 L. Goerigk, Treating London-Dispersion Effects with the Latest Minnesota Density Functionals: Problems and Possible Solutions, J. Phys. Chem. Lett., 2015, 6, 3891–3896, DOI: 10.1021/acs.jpclett.5b01591.
  • Peverati and Truhlar 2012 R. Peverati and D. G. Truhlar, Exchange–Correlation Functional with Good Accuracy for Both Structural and Energetic Properties while Depending Only on the Density and Its Gradient, J. Chem. Theory Comput., 2012, 8, 2310–2319, DOI: 10.1021/ct3002656.
  • Jin and Bartlett 2016 Y. Jin and R. J. Bartlett, The QTP family of consistent functionals and potentials in Kohn–Sham density functional theory, J. Chem. Phys., 2016, 145, 034107, DOI: 10.1063/1.4955497.
  • Perdew et al. 1996 J. P. Perdew, K. Burke and M. Ernzerhof, Generalized Gradient Approximation Made Simple, Phys. Rev. Lett., 1996, 77, 3865–3868, DOI: 10.1103/PhysRevLett.77.3865.
  • Perdew et al. 1997 J. P. Perdew, K. Burke and M. Ernzerhof, Generalized Gradient Approximation Made Simple [Phys. Rev. Lett. 77, 3865 (1996)], Phys. Rev. Lett., 1997, 78, 1396–1396, DOI: 10.1103/PhysRevLett.78.1396.
  • Haiduke and Bartlett 2018 R. L. A. Haiduke and R. J. Bartlett, Non-empirical exchange-correlation parameterizations based on exact conditions from correlated orbital theory, J. Chem. Phys., 2018, 148, 184106, DOI: 10.1063/1.5025723.
  • Zhang and Yang 1998 Y. Zhang and W. Yang, Comment on "Generalized Gradient Approximation Made Simple", Phys. Rev. Lett., 1998, 80, 890–890, DOI: 10.1103/PhysRevLett.80.890.
  • Heyd et al. 2003 J. Heyd, G. E. Scuseria and M. Ernzerhof, Hybrid functionals based on a screened Coulomb potential, J. Chem. Phys., 2003, 118, 8207–8215, DOI: 10.1063/1.1564060.
  • Heyd et al. 2006 J. Heyd, G. E. Scuseria and M. Ernzerhof, Erratum: "Hybrid functionals based on a screened Coulomb potential" [J. Chem. Phys. 118, 8207 (2003)], J. Chem. Phys., 2006, 124, 219906, DOI: 10.1063/1.2204597.
  • Hammer et al. 1999 B. Hammer, L. B. Hansen and J. K. Nørskov, Improved adsorption energetics within density-functional theory using revised Perdew–Burke–Ernzerhof functionals, Phys. Rev. B, 1999, 59, 7413–7421, DOI: 10.1103/PhysRevB.59.7413.
  • Krukau et al. 2006 A. V. Krukau, O. A. Vydrov, A. F. Izmaylov and G. E. Scuseria, Influence of the exchange screening parameter on the performance of screened hybrid functionals, J. Chem. Phys., 2006, 125, 224106, DOI: 10.1063/1.2404663.
  • Reimers et al. 2015 J. R. Reimers, D. Panduwinata, J. Visser, Y. Chin, C. Tang, L. Goerigk, M. J. Ford, M. Sintic, T.-J. Sum, M. J. J. Coenen, B. L. M. Hendriksen, J. A. A. W. Elemans, N. S. Hush and M. J. Crossley, A priori calculations of the free energy of formation from solution of polymorphic self-assembled monolayers, Proc. Natl. Acad. Sci., 2015, 112, E6101, DOI: 10.1073/pnas.1516984112.
  • Perdew et al. 1992 J. P. Perdew, J. A. Chevary, S. H. Vosko, K. A. Jackson, M. R. Pederson, D. J. Singh and C. Fiolhais, Atoms, molecules, solids, and surfaces: Applications of the generalized gradient approximation for exchange and correlation, Phys. Rev. B, 1992, 46, 6671–6687, DOI: 10.1103/PhysRevB.46.6671.
  • Perdew et al. 1993 J. P. Perdew, J. A. Chevary, S. H. Vosko, K. A. Jackson, M. R. Pederson, D. J. Singh and C. Fiolhais, Erratum: Atoms, molecules, solids, and surfaces: Applications of the generalized gradient approximation for exchange and correlation, Phys. Rev. B, 1993, 48, 4978–4978, DOI: 10.1103/PhysRevB.48.4978.2.
  • Peverati and Truhlar 2012 R. Peverati and D. G. Truhlar, Screened-exchange density functionals with broad accuracy for chemistry and solid-state physics, Phys. Chem. Chem. Phys., 2012, 14, 16187–16191, DOI: 10.1039/c2cp42576a.
  • Xu and Goddard 2004 X. Xu and W. A. Goddard, From The Cover: The X3LYP extended density functional for accurate descriptions of nonbond interactions, spin states, and thermochemical properties, Proc. Natl. Acad. Sci. U. S. A., 2004, 101, 2673–2677, DOI: 10.1073/pnas.0308730100.
  • Mardirossian and Head-Gordon 2015 N. Mardirossian and M. Head-Gordon, Mapping the genome of meta-generalized gradient approximation density functionals: The search for B97M-V, J. Chem. Phys., 2015, 142, 074111, DOI: 10.1063/1.4907719.
  • Zhao and Truhlar 2006 Y. Zhao and D. G. Truhlar, A new local density functional for main-group thermochemistry, transition metal bonding, thermochemical kinetics, and noncovalent interactions, J. Chem. Phys., 2006, 125, 194101, DOI: 10.1063/1.2370993.
  • Adamo and Barone 1997 C. Adamo and V. Barone, Toward reliable adiabatic connection models free from adjustable parameters, Chem. Phys. Lett., 1997, 274, 242–250, DOI: 10.1016/S0009-2614(97)00651-9.
  • Wang et al. 2017 Y. Wang, X. Jin, H. S. Yu, D. G. Truhlar and X. He, Revised M06-L functional for improved accuracy on chemical reaction barrier heights, noncovalent interactions, and solid-state physics, Proc. Natl. Acad. Sci. U. S. A., 2017, 114, 8487–8492, DOI: 10.1073/pnas.1705670114.
  • Stephens et al. 1994 P. J. Stephens, F. J. Devlin, C. F. Chabalowski and M. J. Frisch, Ab Initio Calculation of Vibrational Absorption and Circular Dichroism Spectra Using Density Functional Force Fields, J. Phys. Chem., 1994, 98, 11623–11627, DOI: 10.1021/j100096a001.
  • Peverati and Truhlar 2012 R. Peverati and D. G. Truhlar, M11-L: A Local Density Functional That Provides Improved Accuracy for Electronic Structure Calculations in Chemistry and Physics, J. Phys. Chem. Lett., 2012, 3, 117–124, DOI: 10.1021/jz201525m.
  • Lu et al. 2013 L. Lu, H. Hu, H. Hou and B. Wang, An improved B3LYP method in the calculation of organic thermochemistry and reactivity, Comput. Theor. Chem., 2013, 1015, 64–71, DOI: 10.1016/j.comptc.2013.04.009.
  • Peverati and Truhlar 2012 R. Peverati and D. G. Truhlar, An improved and broadly accurate local approximation to the exchange-correlation density functional: The MN12-L functional for electronic structure calculations in chemistry and physics, Phys. Chem. Chem. Phys., 2012, 14, 13171, DOI: 10.1039/c2cp42025b.
  • Becke 1993 A. D. Becke, Density-functional thermochemistry. III. The role of exact exchange, J. Chem. Phys., 1993, 98, 5648–5652, DOI: 10.1063/1.464913.
  • Yu et al. 2016 H. S. Yu, X. He and D. G. Truhlar, MN15-L: A New Local Exchange-Correlation Functional for Kohn–Sham Density Functional Theory with Broad Accuracy for Atoms, Molecules, and Solids, J. Chem. Theory Comput., 2016, 12, 1280–1293, DOI: 10.1021/acs.jctc.5b01082.
  • Sun et al. 2015 J. Sun, J. P. Perdew and A. Ruzsinszky, Semilocal density functional obeying a strongly tightened bound for exchange, Proc. Natl. Acad. Sci. U. S. A., 2015, 112, 685–689, DOI: 10.1073/pnas.1423145112.
  • Perdew et al. 2009 J. P. Perdew, A. Ruzsinszky, G. I. Csonka, L. A. Constantin and J. Sun, Workhorse Semilocal Density Functional for Condensed Matter Physics and Quantum Chemistry, Phys. Rev. Lett., 2009, 103, 026403, DOI: 10.1103/PhysRevLett.103.026403.
  • Becke 1997 A. D. Becke, Density-functional thermochemistry. V. Systematic optimization of exchange-correlation functionals, J. Chem. Phys., 1997, 107, 8554–8560, DOI: 10.1063/1.475007.
  • Sun et al. 2015 J. Sun, A. Ruzsinszky and J. P. Perdew, Strongly Constrained and Appropriately Normed Semilocal Density Functional, Phys. Rev. Lett., 2015, 115, 036402, DOI: 10.1103/PhysRevLett.115.036402.
  • Hamprecht et al. 1998 F. A. Hamprecht, A. J. Cohen, D. J. Tozer and N. C. Handy, Development and assessment of new exchange-correlation functionals, J. Chem. Phys., 1998, 109, 6264, DOI: 10.1063/1.477267.
  • Bartók and Yates 2019 A. P. Bartók and J. R. Yates, Regularized SCAN functional, J. Chem. Phys., 2019, 150, 161101, DOI: 10.1063/1.5094646.
  • Wilson et al. 2001 P. J. Wilson, T. J. Bradley and D. J. Tozer, Hybrid exchange-correlation functional determined from thermochemical data and ab initio potentials, J. Chem. Phys., 2001, 115, 9233, DOI: 10.1063/1.1412605.
  • Ehlert et al. 2021 S. Ehlert, U. Huniar, J. Ning, J. W. Furness, J. Sun, A. D. Kaplan, J. P. Perdew and J. G. Brandenburg, r2SCAN-D4: Dispersion corrected meta-generalized gradient approximation for general chemical applications, J. Chem. Phys., 2021, 154, 061101, DOI: 10.1063/5.0041008.
  • Furness et al. 2020 J. W. Furness, A. D. Kaplan, J. Ning, J. P. Perdew and J. Sun, Accurate and Numerically Efficient r2SCAN Meta-Generalized Gradient Approximation, J. Phys. Chem. Lett., 2020, 11, 8208–8215, DOI: 10.1021/acs.jpclett.0c02405.
  • Furness et al. 2020 J. W. Furness, A. D. Kaplan, J. Ning, J. P. Perdew and J. Sun, Correction to "Accurate and Numerically Efficient r2SCAN Meta-Generalized Gradient Approximation", J. Phys. Chem. Lett., 2020, 11, 9248–9248, DOI: 10.1021/acs.jpclett.0c03077.
  • Keal and Tozer 2005 T. W. Keal and D. J. Tozer, Semiempirical hybrid functional with improved performance in an extensive chemical assessment, J. Chem. Phys., 2005, 123, 121103, DOI: 10.1063/1.2061227.
  • Aschebrock and Kümmel 2019 T. Aschebrock and S. Kümmel, Ultranonlocality and accurate band gaps from a meta-generalized gradient approximation, Phys. Rev. Res., 2019, 1, 033082, DOI: 10.1103/PhysRevResearch.1.033082.
  • Perdew and Wang 1992 J. P. Perdew and Y. Wang, Accurate and simple analytic representation of the electron-gas correlation energy, Phys. Rev. B, 1992, 45, 13244–13249, DOI: 10.1103/PhysRevB.45.13244.
  • Becke 1993 A. D. Becke, A new mixing of Hartree–Fock and local density-functional theories, J. Chem. Phys., 1993, 98, 1372–1377, DOI: 10.1063/1.464304.
  • Tao et al. 2003 J. Tao, J. P. Perdew, V. N. Staroverov and G. E. Scuseria, Climbing the Density Functional Ladder: Nonempirical Meta-Generalized Gradient Approximation Designed for Molecules and Solids, Phys. Rev. Lett., 2003, 91, 146401, DOI: 10.1103/PhysRevLett.91.146401.
  • Perdew et al. 2004 J. P. Perdew, J. Tao, V. N. Staroverov and G. E. Scuseria, Meta-generalized gradient approximation: Explanation of a realistic nonempirical density functional, J. Chem. Phys., 2004, 120, 6898–6911, DOI: 10.1063/1.1665298.
  • Kruse et al. 2018 H. Kruse, P. Banáš and J. Šponer, Investigations of Stacked DNA Base-Pair Steps: Highly Accurate Stacking Interaction Energies, Energy Decomposition, and Many-Body Stacking Effects, J. Chem. Theory Comput., 2018, 15, 95–115, DOI: 10.1021/acs.jctc.8b00643.
  • Perdew et al. 2011 J. P. Perdew, A. Ruzsinszky, G. I. Csonka, L. A. Constantin and J. Sun, Erratum: Workhorse Semilocal Density Functional for Condensed Matter Physics and Quantum Chemistry [Phys. Rev. Lett. 103, 026403 (2009)], Phys. Rev. Lett., 2011, 106, 179902, DOI: 10.1103/PhysRevLett.106.179902.
  • Boese and Martin 2004 A. D. Boese and J. M. L. Martin, Development of density functionals for thermochemical kinetics, J. Chem. Phys., 2004, 121, 3405, DOI: 10.1063/1.1774975.
  • Boese et al. 2000 A. D. Boese, N. L. Doltsinis, N. C. Handy and M. Sprik, New generalized gradient approximation functionals, J. Chem. Phys., 2000, 112, 1670, DOI: 10.1063/1.480732.
  • Wang et al. 2018 Y. Wang, P. Verma, X. Jin, D. G. Truhlar and X. He, Revised M06 density functional for main-group and transition-metal chemistry, Proc. Natl. Acad. Sci. U. S. A., 2018, 115, 10257–10262, DOI: 10.1073/pnas.1810421115.
  • Boese and Handy 2001 A. D. Boese and N. C. Handy, A new parametrization of exchange–correlation generalized gradient approximation functionals, J. Chem. Phys., 2001, 114, 5497, DOI: 10.1063/1.1347371.
  • Adamo and Barone 1998 C. Adamo and V. Barone, Exchange functionals with improved long-range behavior and adiabatic connection methods without adjustable parameters: The mPW and mPW1PW models, J. Chem. Phys., 1998, 108, 664–675, DOI: 10.1063/1.475428.
  • Zhao and Truhlar 2008 Y. Zhao and D. G. Truhlar, Exploring the Limit of Accuracy of the Global Hybrid Meta Density Functional for Main-Group Thermochemistry, Kinetics, and Noncovalent Interactions, J. Chem. Theory Comput., 2008, 4, 1849–1868, DOI: 10.1021/ct800246v.
  • Hoe et al. 2001 W.-M. Hoe, A. J. Cohen and N. C. Handy, Assessment of a new local exchange functional OPTX, Chem. Phys. Lett., 2001, 341, 319–328, DOI: 10.1016/S0009-2614(01)00581-4.
  • Cohen and Handy 2001 A. J. Cohen and N. C. Handy, Dynamic correlation, Mol. Phys., 2001, 99, 607–615, DOI: 10.1080/00268970010023435.
  • Adamo and Barone 1999 C. Adamo and V. Barone, Toward reliable density functional methods without adjustable parameters: The PBE0 model, J. Chem. Phys., 1999, 110, 6158–6170, DOI: 10.1063/1.478522.
  • Ernzerhof and Scuseria 1999 M. Ernzerhof and G. E. Scuseria, Assessment of the Perdew–Burke–Ernzerhof exchange-correlation functional, J. Chem. Phys., 1999, 110, 5029–5036, DOI: 10.1063/1.478401.
  • Yu et al. 2016 H. S. Yu, X. He, S. L. Li and D. G. Truhlar, MN15: A Kohn–Sham global-hybrid exchange–correlation density functional with broad accuracy for multi-reference and single-reference systems and noncovalent interactions, Chem. Sci., 2016, 7, 5032–5051, DOI: 10.1039/C6SC00705H.
  • Jin and Bartlett 2018 Y. Jin and R. J. Bartlett, Accurate computation of X-ray absorption spectra with ionization potential optimized global hybrid functional, J. Chem. Phys., 2018, 149, 064111, DOI: 10.1063/1.5038434.
  • Csonka et al. 2010 G. I. Csonka, J. P. Perdew and A. Ruzsinszky, Global Hybrid Functionals: A Look at the Engine under the Hood, J. Chem. Theory Comput., 2010, 6, 3688–3703, DOI: 10.1021/ct100488v.
  • Peverati and Truhlar 2011 R. Peverati and D. G. Truhlar, Communication: A global hybrid generalized gradient approximation to the exchange-correlation functional that satisfies the second-order density-gradient constraint and has broad applicability in chemistry, J. Chem. Phys., 2011, 135, 191102, DOI: 10.1063/1.3663871.
  • Mardirossian and Head-Gordon 2016 N. Mardirossian and M. Head-Gordon, $\omega$B97M-V: A combinatorially optimized, range-separated hybrid, meta-GGA density functional with VV10 nonlocal correlation, J. Chem. Phys., 2016, 144, 214110, DOI: 10.1063/1.4952647.
  • Wang et al. 2020 Y. Wang, P. Verma, L. Zhang, Y. Li, Z. Liu, D. G. Truhlar and X. He, M06-SX screened-exchange density functional for chemistry and solid-state physics, Proc. Natl. Acad. Sci. U. S. A., 2020, 117, 2294–2301, DOI: 10.1073/pnas.1913699117.
  • Peverati and Truhlar 2011 R. Peverati and D. G. Truhlar, Improving the Accuracy of Hybrid Meta-GGA Density Functionals by Range Separation, J. Phys. Chem. Lett., 2011, 2, 2810–2817, DOI: 10.1021/jz201170d.
  • Verma et al. 2019 P. Verma, Y. Wang, S. Ghosh, X. He and D. G. Truhlar, Revised M11 Exchange-Correlation Functional for Electronic Excitation Energies and Ground-State Properties, J. Phys. Chem. A, 2019, 123, 2966–2990, DOI: 10.1021/acs.jpca.8b11499.
  • Johnson and Becke 2005 E. R. Johnson and A. D. Becke, A post-Hartree–Fock model of intermolecular interactions, J. Chem. Phys., 2005, 123, 024101, DOI: 10.1063/1.1949201.
  • Lin et al. 2013 Y.-S. Lin, G.-D. Li, S.-P. Mao and J.-D. Chai, Long-Range Corrected Hybrid Density Functionals with Improved Dispersion Corrections, J. Chem. Theory Comput., 2013, 9, 263–272, DOI: 10.1021/ct300715s.
Graphical abstract
Hydrogen-bonding strengths of quadruply bonded self-complementary dimers are calculated using 152 functionals and compared to coupled-cluster reference energies extrapolated to the Schrödinger limit.

### v2

@@
 In this work, we study the reproduction of these bonding energies at the DFT level using 152 DFAs.
-The top ten functionals are composed of eight variants of the Berkeley functionals both with and without dispersion corrections, and two Minnesota 2011 functionals augmented with a further dispersion correction.
-We find the B97M-V functional with the non-local correlation functional replaced by an empirical D3BJ dispersion correction to be the best functional, while changes to the dispersion part in other Berkeley functionals lead to poorer performance in our study.
+The top ten density functional approximations are composed of eight variants of the Berkeley functionals both with and without dispersion corrections, and two Minnesota 2011 functionals augmented with a further dispersion correction.
+We find the B97M-V functional with the non-local correlation functional replaced by an empirical D3BJ dispersion correction to be the best DFA, while changes to the dispersion part in other Berkeley functionals lead to poorer performance in our study.
 1 Introduction
@@
 The accuracy of various DFAs for modeling hydrogen bonds has been evaluated in a number of works.25, 26, 27, 28, 29 More recent DFA benchmark studies of hydrogen-bonding strengths have employed the S22, S66 and S66x8 datasets.30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45 The pronounced relevance of hydrogen bonding in chemistry and their DFT modeling is underlined by the fact that hydrogen bond strengths have been included in the fitting of the Minnesota functionals since their first versions,29, 46, 47 for example.
-In this work, we study how well various DFAs reproduce the hydrogen bonding in 14 quadruply hydrogen-bonded dimers.
-Half of the molecules exhibit the DDAA-AADD and the other half exhibits the DADA-ADAD hydrogen bonding motif.
-We consider 152 DFAs including the Hartree–Fock (HF) method.
-The employed dispersion correction schemes include Grimme’s empirical dispersion corrections (DFT-D348, DFT-D449) and a non-local (NL) correlation functional (VV10).50
+[…]
 Some of us have recently reported accurate reference values for the molecules studied in this work, which were calculated with domain-based local pair natural orbital coupled-cluster theory at the singles, doubles and perturbative triples DLPNO-CCSD(T) level extrapolated to the complete basis set limit.1 Ahmed et al.1 also extrapolated the reference energies towards the Schrödinger limit following Goodson51, producing the data that we will use as reference in this work.
@@
 2 Computational methods
-The benchmark considers the 14 molecules of the study of Ahmed et al.1 that determined hydrogen-bonding energies for TPSSh-D3/def2-TZVPP optimized geometries at the DLPNO-CCSD(T) level of theory extrapolated to the complete basis set limit, as well as towards the Schrödinger limit using the continued-fraction (cf) scheme of Goodson51.
+The benchmark considers the 14 dimers of the study of Ahmed et al.1 that determined hydrogen-bonding energies for TPSSh-D3/def2-TZVPP optimized geometries at the DLPNO-CCSD(T) level of theory extrapolated to the complete basis set limit, as well as towards the Schrödinger limit using the continued-fraction (cf) scheme of Goodson51.
 We denote the reference level of theory as CCSD(T)-cf.
-The reference energies are taken from Ahmed et al.1 and are included in the ESI†along with the employed geometries and visualizations thereof.
+The reference energies and molecular structures of the dimers reported in the ESI† are taken from Ahmed et al.1.
+The molecular structures of the monomers with the DADA and DDAA motifs are shown in Figs.
+1 and 2, respectively.
+(a) DADA01
+(b) DADA02
+(c) DADA03
+(d) DADA04
+(e) DADA05
+(f) DADA06
+(g) DADA08
+Fig.
+1: The monomers of the quadruple hydrogen-bonded dimers with DADA-ADAD hydrogen-bonding pattern.
+The previously studied DADA071 is omitted because the CCSD(T)-cf reference data are not available.
+The figures are made with Jmol52.
+(a) DDAA02
+(b) DDAA03
+(c) DDAA04
+(d) DDAA05
+(e) DDAA06
+(f) DDAA07
+(g) DDAA08
+Fig.
+2: The monomers of the quadruple hydrogen-bonded dimers with the DDAA-AADD hydrogen-bonding pattern.
+The previously studied DDAA011 is omitted because the CCSD(T)-cf reference data are not available.
+The figures are made with Jmol52.
 […]
-The calculations were performed using the default (75,302) exchange-correlation grid in Psi4, which uses 75 radial grid points placed following Treutler and Ahlrichs55, and an angular Lebedev grid with 302 points.56 The VV10 non-local correlation functional50 was likewise evaluated on the corresponding (50,146) default grid.
-The resolution-of-the-identity approximation was used in all calculations with the universal Hartree–Fock auxiliary basis sets.57
+The hydrogen-bonding energies were corrected for basis set superposition errors (BSSE) calculated using the counterpoise correction method56, 57.
+The BSSE was obtained by comparing the energies of the two monomers of the dimer structure calculated using the dimer basis set to the monomer energies calculated using the monomer basis sets and the same molecular structure.
+The BSSE was added to the hydrogen-bonding energy calculated for fully optimized monomers using the basis set of the monomers.
+The calculations were performed using the default (75,302) exchange-correlation grid in Psi4, which uses 75 radial grid points placed following Treutler and Ahlrichs58, and an angular Lebedev grid with 302 points.59 The VV10 non-local correlation functional50 was likewise evaluated on the corresponding (50,146) default grid.
+The resolution-of-the-identity approximation was used in all calculations with the universal Hartree–Fock auxiliary basis sets.60
 The 75 base DFAs considered in this work, which have been listed in Table 1, comprise 14 semi-local generalized gradient approximation (GGA) functionals, 21 global hybrid (GH) GGA functionals, 13 range-separated (RS) GGA functionals, 13 semi-local meta-GGA (mGGA) functionals, 8 GH mGGA functionals, and 5 RS mGGA functionals, as well as HF (a GH functional without a semi-local exchange-correlation part).
@@
 Versions 6.0.0 and 6.2.2 of the library were employed.
-We note that some functionals such as B3LYP, HSE03, and HSE06 have various definitions in the literature,58 which is avoided through the use of the standard implementations in Libxc.
+We note that some functionals such as B3LYP, HSE03, and HSE06 have various definitions in the literature,61 which is avoided through the use of the standard implementations in Libxc.
 Table 1 also shows the empirical (DFT-D348, DFT-D449) dispersion corrections or non-local correlation (VV1050) additionally considered with some of the functionals.
 Various combinations of these dispersion models (or the lack thereof) with the 75 base DFAs yield the 152 DFAs investigated in this study.
+The study by Ahmed et al.1 showed that solvent effects increase the hydrogen-bonding energy of the tautomers DADA08 and DDAA06 by only 4-5%.
+We have not considered any solvent effects in this work because they are relatively small.
+Solvent effects are not expected to significantly affect the ranking of the DFAs.
 Table 1: The density functional approximations considered in this work. The three dispersion correction schemes are also indicated.
 functional type D3 D4 NL ref. functional type D3 D4 NL ref.
-B97-D3 GGA 59a 59 $\omega$B97 RS GGA 60
-BLYP GGA 48a 49 61 62, 63, 64 $\omega$B97X RS GGA 49 60
-BOP GGA 48a 65 $\omega$B97X-Db RS GGA 66
-BP86 GGA 35a 67 62, 68 $\omega$B97X-V RS GGA 69c 70
-CHACHIYO GGA 71, 72 $\omega$PBE RS GGA 48a 73, 74
-KT1 GGA 75 CAM-B3LYP RS GGA 48a 76
-KT2 GGA 75 CAMh-B3LYP RS GGA 77
-KT3 GGA 78 CAM-QTP-00 RS GGA 79
-N12 GGA 80a 81 CAM-QTP-01 RS GGA 82
-PBE GGA 48a 49 67 83, 84 CAM-QTP-02 RS GGA 85
-revPBE GGA 48a 49 61 86 HSE03 RS GGA 48a 49 87, 88
-rPBE GGA 48a 49 89 HSE06 RS GGA 48a 49 87, 88, 90
-PW91 GGA 91a 49 92, 93 N12-SX RS GGA 80a 94
-XLYP GGA 48a 49 95 B97M-V mGGA 69c 96
-HF GH 48a 49 M06-L mGGA 97
-B1LYP GH GGA 48a 49 98 revM06-L mGGA 99
-B3LYP GH GGA 48a 49 61 100 M11-L mGGA 80a 101
-revB3LYP GH GGA 102 MN12-L mGGA 80a 103
-B3PW91 GH GGA 48a 61 104 MN15-L mGGA 105
-B3P86 GH GGA 35a 104 MVS mGGA 106, 107
-B97 GH GGA 48c 108 SCAN mGGA 48a 49 109
-B97-1 GH GGA 35a 110 rSCAN mGGA 111
-B97-2 GH GGA 35a 112 r2SCAN mGGA 48a 113 114, 115
-B97-3 GH GGA 116 TASK mGGA 117, 118
-BHandH GH GGA 119 TPSS mGGA 48a 49 61 120, 121
-BHandHLYP GH GGA 62, 119 revTPSS mGGA 48a 49 122 107, 123
-BMK GH GGA 48a 124 M06 GH mGGA 31
-HCTH120 GH GGA 48a 125 revM06 GH mGGA 126
-HCTH407 GH GGA 48a 127 M06-2X GH mGGA 31
-MPW1PW GH GGA 48a 49 128 M08-HX GH mGGA 129
-O3LYP GH GGA 48a 49 130, 131 M08-SO GH mGGA 129
-PBE0 GH GGA 48a 49 67 132, 133 MN15 GH mGGA 48a 134
-revPBE0 GH GGA 48a 49 61 86, 83, 84 TPSSh GH mGGA 48a 49 27
-QTP17 GH GGA 135 revTPSSh GH mGGA 48a 49 136
-SOGGA11-X GH GGA 80a 137 $\omega$B97M-V RS mGGA 69c 138
-X3LYP GH GGA 48a 49 95 M06-SX RS mGGA 139
-M11 RS mGGA 48a 140
-revM11 RS mGGA 141
-MN12-SX RS mGGA 80a 94
+B97-D3 GGA 62a 62 $\omega$B97 RS GGA 63
+BLYP GGA 48a 49 64 65, 66, 67 $\omega$B97X RS GGA 49 63
+BOP GGA 48a 68 $\omega$B97X-Db RS GGA 69
+BP86 GGA 35a 70 65, 71 $\omega$B97X-V RS GGA 72c 73
+CHACHIYO GGA 74, 75 $\omega$PBE RS GGA 48a 76, 77
+KT1 GGA 78 CAM-B3LYP RS GGA 48a 79
+KT2 GGA 78 CAMh-B3LYP RS GGA 80
+KT3 GGA 81 CAM-QTP-00 RS GGA 82
+N12 GGA 83a 84 CAM-QTP-01 RS GGA 85
+PBE GGA 48a 49 70 86, 87 CAM-QTP-02 RS GGA 88
+revPBE GGA 48a 49 64 89 HSE03 RS GGA 48a 49 90, 91
+rPBE GGA 48a 49 92 HSE06 RS GGA 48a 49 90, 91, 93
+PW91 GGA 94a 49 95, 96 N12-SX RS GGA 83a 97
+XLYP GGA 48a 49 98 B97M-V mGGA 72c 99
+HF GH 48a 49 M06-L mGGA 100
+B1LYP GH GGA 48a 49 101 revM06-L mGGA 102
+B3LYP GH GGA 48a 49 64 103 M11-L mGGA 83a 104
+revB3LYP GH GGA 105 MN12-L mGGA 83a 106
+B3PW91 GH GGA 48a 64 107 MN15-L mGGA 108
+B3P86 GH GGA 35a 107 MVS mGGA 109, 110
+B97 GH GGA 48c 111 SCAN mGGA 48a 49 112
+B97-1 GH GGA 35a 113 rSCAN mGGA 114
+B97-2 GH GGA 35a 115 r2SCAN mGGA 48a 116 117, 118
+B97-3 GH GGA 119 TASK mGGA 120, 121
+BHandH GH GGA 122 TPSS mGGA 48a 49 64 123, 124
+BHandHLYP GH GGA 65, 122 revTPSS mGGA 48a 49 125 110, 126
+BMK GH GGA 48a 127 M06 GH mGGA 31
+HCTH120 GH GGA 48a 128 revM06 GH mGGA 129
+HCTH407 GH GGA 48a 130 M06-2X GH mGGA 31
+MPW1PW GH GGA 48a 49 131 M08-HX GH mGGA 132
+O3LYP GH GGA 48a 49 133, 134 M08-SO GH mGGA 132
+PBE0 GH GGA 48a 49 70 135, 136 MN15 GH mGGA 48a 137
+revPBE0 GH GGA 48a 49 64 89, 86, 87 TPSSh GH mGGA 48a 49 27
+QTP17 GH GGA 138 revTPSSh GH mGGA 48a 49 139
+SOGGA11-X GH GGA 83a 140 $\omega$B97M-V RS mGGA 72c 141
+X3LYP GH GGA 48a 49 98 M06-SX RS mGGA 142
+M11 RS mGGA 48a 143
+revM11 RS mGGA 144
+MN12-SX RS mGGA 83a 97
   • a
-    Empirical dispersion with Becke–Johnson142 3-body damping.
+    Empirical dispersion with Becke–Johnson145 3-body damping.
   • b
@@
   • c
-    Empirical dispersion with Becke–Johnson142 2-body damping.
+    Empirical dispersion with Becke–Johnson145 2-body damping.
 3 Results
@@
 The range of truncation errors for the polarized split-valence basis sets drops significantly from $\approx$ 0.8–7 kJ/mol with def2-SVP in Fig.
-1(a) to $\approx$ $-1.5$–0.5 kJ/mol with def2-SVPD; this highlights the need to include diffuse functions for describing non-covalent bonding, as in the present hydrogen bonded systems.
+3(a) to $\approx$ $-1.5$–0.5 kJ/mol with def2-SVPD; this highlights the need to include diffuse functions for describing non-covalent bonding, as in the present hydrogen bonded systems.
 A convergence to within ±$\pm 0.5$ kJ/mol precision with respect to the basis set is already achieved with the def2-TZVPD basis set.
@@
 (b) Convergence with diffuse functions
-Fig.  1: Basis-set convergence plot of the hydrogen-bonding energies, relative to B97M-D3/def2-QZVPPD values. The figures show an energy window of ±$\pm 0.5$ kJ/mol as horizontal dashed lines.
-3.2 The best functionals
-The functionals were ranked based on their mean absolute errors (MAE).
+Fig.  3: Basis-set convergence plot of the hydrogen-bonding energies, relative to B97M-D3/def2-QZVPPD values. The figures show an energy window of ±$\pm 0.5$ kJ/mol as horizontal dashed lines.
+3.2 The best density functional approximations
+The density functional approximations were ranked based on their mean absolute errors (MAE).
 The ranking is shown in Tables 2 and 3 together with the corresponding MAEs and maximum absolute errors (MAX).
@@
 Replacing VV10 with the D3 dispersion correction drops the ranking to the 7th place.
-Table 2: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the 76 highest ranked functionals.
-The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The functionals are ordered in increasing MAE.
+Table 2: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the 76 highest ranked density functional approximations.
+The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The DFAs are ordered in increasing MAE.
 Rank Functional MAE MAX Rank Functional MAE MAX
@@
 38 revPBE0-D4 6.01 11.03 76 TPSS-NL 10.88 16.88
-Table 3: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the functionals ranked from 77 to 152.
-The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The functionals are ordered in increasing MAE.
+Table 3: The mean absolute errors (MAE) and maximum absolute errors (MAX) for the hydrogen-bonding strength of the 14 studied molecules (in kJ/mol) using the density functional approximations ranked from 77 to 152.
+The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1 The DFAs are ordered in increasing MAE.
 Rank Functional MAE MAX Rank Functional MAE MAX
@@
 The Minnesota M11-L-D3 holds 3rd place on the list with the added D3 dispersion corrections, while the original functional parametrized without the empirical dispersion term (M11-L) holds the 104th place.
-The functionals ranked $5^{\mathrm{th}}$ to $9^{\mathrm{th}}$ are all Berkeley $\omega$B97 (RS GGA) functionals with different dispersion correction terms.
-$\omega$B97X-D3 is the $\omega$B97X-V70 functional with a D3 Becke–Johnson142, 48 dispersion correction instead of VV1069, and it should not to be confused with the $\omega$B97X-D3 functional of Lin et al.143 which was not studied in this work.
+The DFAs ranked $5^{\mathrm{th}}$ to $9^{\mathrm{th}}$ are all Berkeley $\omega$B97 (RS GGA) functionals with different dispersion correction terms.
+$\omega$B97X-D3 is the $\omega$B97X-V73 functional with a D3 Becke–Johnson145, 48 dispersion correction instead of VV1072, and it should not to be confused with the $\omega$B97X-D3 functional of Lin et al.146 which was not studied in this work.
 The Minnesota M11 functional with an additional D3 correction is on the 10th place, while the original functional without the dispersion term ranks 65th.
@@
 mGGA and global hybrid (GH) GGA functionals also appear among the 20 most accurate functionals for hydrogen-bonding energies.
-The best GGA functionals are ranked 17th (XLYP-D3), 19th (rPBE-D4), 21st (BLYP-D3) and 22nd (XLYP-D4).
+The best DFAs at the GGA level are ranked 17th (XLYP-D3), 19th (rPBE-D4), 21st (BLYP-D3) and 22nd (XLYP-D4).
 Fig.
-2: Violin plots showing the error distributions in the hydrogen-bonding energies (in kJ/mol) for the investigated DFAs.
+4: Violin plots showing the error distributions in the hydrogen-bonding energies (in kJ/mol) for the investigated DFAs.
 The reference hydrogen-bonding energies are extrapolated CCSD(T)-cf values from the literature.1
-The error distributions for the hydrogen-bonding energies of the 36 top functionals are shown as violin plots in Fig.
-2, while violin plots for the remaining functionals are included in the supporting information (SI).
-The average errors in the hydrogen-bonding energies for the top functionals are close to zero, except for M11-D3, whose distribution is almost completely on the negative side, implying that it systematically underestimates the hydrogen bonding energies.
-Similar plots are shown for the functionals ranked 13th–24th in Fig.
-2, where the distributions are broader, and the average is larger than zero for many functionals, which implies that the hydrogen-bonding energies are overestimated.
-The error distributions for the functionals ranked between 25th and 36th shown in Fig.
-2 are broader and more scattered than for the higher-ranked functionals.
-The violin plots for the rest of the functionals are given in the ESI†.
+The error distributions for the hydrogen-bonding energies of the 36 top DFAs are shown as violin plots in Fig.
+4, while violin plots for the remaining DFAs are included in the ESI†.
+The average errors in the hydrogen-bonding energies for the top DFAs are close to zero, except for M11-D3, whose distribution is almost completely on the negative side, implying that it systematically underestimates the hydrogen bonding energies.
+Similar plots are shown for the DFAs ranked 13th–24th in Fig.
+4, where the distributions are broader, and the average is larger than zero for many DFAs, which implies that the hydrogen-bonding energies are overestimated.
+The error distributions for the DFAs ranked between 25th and 36th shown in Fig.
+4 are broader and more scattered than for the higher-ranked DFAs.
+The violin plots for the rest of the DFAs are given in the ESI†.
+Hydrogen-bonding energies have been previously calculated in the basis-set limit at the second-order Møller-Plesset perturbation theory (MP2) level20.
+Since all deviations have the same sign, MP2 systematically overestimates the hydrogen-bonding energies, whereas the deviations of the hydrogen-bonding energies calculated using the top DFAs are both positive and negative.
+The hydrogen-bonding energies at the MP2 level correlate linearly with the reference energies.
+The angular coefficient of the linear fit of the MP2 energies is $1.022\pm 0.02$ with an offset of $-5.4\pm 3.3$ kJ/mol.
+The MAE and MAX values at the MP2 level are 8.3 kJ/mol and 14.6 kJ/mol, respectively, corresponding to the 55th place in the ranking list.
 3.3 Meta-GGA functionals
-The best performing mGGA functionals are B97M-D3 (1st), M11-L-D3 (3rd) and B97M-V (4th).
+The best performing DFAs at the mGGA level are B97M-D3 (1st), M11-L-D3 (3rd) and B97M-V (4th).
 SCAN and revTPSS-D3 are mGGA functionals that also perform well, holding rank 12 and 13, respectively.
@@
 The best of the rest of the RS mGGA functionals is revM11, which is ranked 50th.
-Three of the five RS GGA functionals from Berkeley are ranked among the 10 best functionals, while adding a dispersion correction to $\omega$B97X leads to worse performace and 42nd rank.
+Three of the five RS GGA functionals from Berkeley are ranked among the 10 best DFAs, while adding a dispersion correction to $\omega$B97X leads to worse performace and 42nd rank.
 The $\omega$B97X-D functional performs poorly in our study (119th), but this is likely caused by lack of dispersion in our calculations with this functional; see Table 1.
@@
 Adding an empirical dispersion correction to the revTPSSh functional improves its performance significantly: the ranking of revTPSSh increases from 125th to 11th (revTPSSh-D3) or 16th (revTPSSh-D4).
+MAX values generally increase with increasing MAE value.
+However, revPBE0-D3 and M06-2X have much smaller MAX values than some of the higher ranked DFAs.
+Judged from the MAX value, revPBE0-D3 and M06-2X would be ranked 12th and 17th instead of 28th and 29th based on the MAE value, respectively.
 3.6 GGA functionals
-The revPBE-NL and XLYP-D3 functionals are the best performing functionals among the GGAs.
+The revPBE-NL and XLYP-D3 functionals are the best performing DFAs among the GGAs.
 Their performance drops from the 14th place to 124th, and 17th rank to 141th without dispersion corrections, respectively.
@@
 4 Summary and conclusions
-We calculated the hydrogen-bonding strength for 14 molecules using 152 density functional approximations (DFAs).
+We calculated the hydrogen-bonding strength for 14 dimers using 152 density functional approximations (DFAs).
 The accuracy of the calculated bond strengths with various DFAs was assessed by comparison to the literature CCSD(T)-cf energies of Ahmed et al.1, which they extrapolated to the complete basis set limit, as well as to the Schrödinger limit using the continued-fraction methodology of Goodson51.
@@
 As the def2-TZVPD and def2-TZVPPD values are within 0.5 kJ/mol of the def2-QZVPPD data, we can safely assume that the quadruple-$\zeta$ reference calculations are indeed of benchmark quality.
-The def2-QZVPPD basis set is often used for benchmarking in the literature, and has been employed in the fitting of the $\omega$B97M-V functional,138 for example.
+The def2-QZVPPD basis set is often used for benchmarking in the literature, and has been employed in the fitting of the $\omega$B97M-V functional,141 for example.
 Having validated the numerical methodology, we continued with the benchmark.
 We found the top of the list to be dominated by the Berkeley functionals, either with or without dispersion.
-The most accurate functional for hydrogen-bonding strengths was found to be B97M-D3, which is the B97M-V functional whose VV10 dispersion correction has been replaced by the empirical D3 dispersion correction.
+The most accurate DFA for hydrogen-bonding strengths was found to be B97M-D3, which is the B97M-V functional whose VV10 dispersion correction has been replaced by the empirical D3 dispersion correction.
 We found the accuracy of mGGA functionals to be surprisingly sensitive to the description of the dispersion energy, tentatively supporting the recent findings of Montero de Hijes et al.24.
-Other functionals that accurately predicted hydrogen-bonding strengths included the range-separated $\omega$B97M-V mGGA functional, the $\omega$B97, $\omega$B97X, $\omega$B97X-V range-separated GGA functionals, as well as the semi-local M11-L-D3 mGGA and the range-separated M11-D3 mGGA functional.
+Other DFAs that accurately predicted hydrogen-bonding strengths included the range-separated $\omega$B97M-V mGGA functional, the $\omega$B97, $\omega$B97X, $\omega$B97X-V range-separated GGA functionals, as well as the semi-local M11-L-D3 mGGA and the range-separated M11-D3 mGGA functional.
 Acknowledgement
@@
   • Goodson 2002 D. Z. Goodson, Extrapolating the coupled-cluster sequence toward the full configuration-interaction limit, J. Chem. Phys., 2002, 116, 6948–6956, DOI: 10.1063/1.1462620.
+  • 52 Jmol: an open-source Java viewer for chemical structures in 3D. http://www.jmol.org.
   • Smith et al.
@@
   • Rappoport and Furche 2010 D. Rappoport and F. Furche, Property-optimized Gaussian basis sets for molecular response calculations, J. Chem. Phys., 2010, 133, 134105, DOI: 10.1063/1.3484283.
+  • Jansen and Ros 1969 H.
+Jansen and P.
+Ros, Non-empirical molecular orbital calculations on the protonation of carbon monoxide, Chem.
+Phys.
+Lett., 1969, 3, 140–143, DOI: 10.1016/0009-2614(69)80118-1.
+  • Boys and Bernardi 1970 S.
+F.
+Boys and F.
+Bernardi, The calculation of small molecular interactions by the differences of separate total energies.
+Some procedures with reduced errors, Mol.
+Phys., 1970, 19, 553–566, DOI: 10.1080/00268977000101561.
   • Treutler and Ahlrichs 1995 O. Treutler and R. Ahlrichs, Efficient molecular numerical integration schemes, J. Chem. Phys., 1995, 102, 346–354, DOI: 10.1063/1.469408.